write an equation for the line with slope 2/3 and y-intercept (0,5)

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

y=mx+b
y=2/3x+5

Answer by Alan3354(69443)

y = mx + b is the slope-intercept form, where
m is the slope and b is the y-intercept.
--------
y = (2/3)x + 5
That's all it takes.